OpenStudy (anonymous):

use the formula to evaluate the series 3+6+12+24+48+96

OpenStudy (paxpolaris):

\[\Large \sum_{k=1}^{6}3 \cdot \left( 2 \right)^{k-1}\]

OpenStudy (anonymous):

ty

OpenStudy (paxpolaris):

\[\Large =3\times {2^6-1 \over 2-1}\]

OpenStudy (anonymous):

thanks